HETAV PATEL

Grade 11,

If Sin (270 0 – X) = COS 292 0 then find x in (0, 360 0 )

If Sin (270 0 – X) = COS 292 0 then find x in (0, 360 0 )

Grade:12th pass

1 Answers

BALAJI ANDALAMALA
8 years ago
sin(270-x) = -cosx
cos(270+22) = cos 22

-cosx = cos 22
cos(180-x) = cos 22
180-x = 22
\therefore x = 158^0